# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v3

@@ 1 Introduction
Learning forms the foundation for intelligent systems to adapt to dynamic environments in response to external changes. Unlike conventional deep neural networks (DNNs) that are designed for static data distribution, Continual Learning (CL) is characterized by learning from sequential, dynamic data distribution [wang2023comprehensive]. However, CL has remained a significant challenge for DNNs due to the issue of catastrophic forgetting: A phenomenon in which adaptation to new tasks results in significant deterioration of consolidated knowledge [masana2022class, de2021continual, ditzler2015learning, robins1995, pmlr-v199-bhat22a]. A number of approaches such as rehearsal-based (e.g., [buzzega2020dark, bhat2022consistency]), weight-regularization based (e.g., [kirkpatrick2017overcoming]) and parameter-isolation based (e.g., [rusu2016progressive]) have been proposed in the literature to address the problem of catastrophic forgetting in CL. Although CL is primarily intended to ensure the resource efficiency of model updates [wang2023comprehensive], the reduction in catastrophic forgetting in these approaches is reliant on increases in computational and memory footprint due to model surrogates and memory buffer. Extrapolating forward this reliance in real world scenarios is both economically and environmentally infeasible.
@@
The increased availability of pre-trained models that demonstrate strong generalization capabilities has led researchers to discover early exploratory methods [douillard2022dytox, jeeveswaran2023birt, gao2024consistent, yang2024generating], facilitating more efficient integration of new knowledge in CL systems. However, continuous fine-tuning on multiple downstream tasks compromises pre-trained model’s generalization capabilities [liu2024parameter]. Recently, Parameter Efficient Fine-Tuning (PEFT) CL techniques [liang2024inflora, zhou2024expandable] are on the rise as they provide lightweight training schedule while avoiding degradation of consolidated knowledge in pre-trained models. However, it is not clear how one can select the size of their rank (in case of LoRA [hu2021lora]) while their performance is very sensitive to rank selection [valipour2022dylora]. Although choosing a higher rank usually results in optimal performance, it significantly increases the number of learnable parameters in longer task sequences when using large pre-trained models. Task arithmetic [ilharcoediting], which enables model compression using simple arithmetic operations with task vectors (difference between pre-trained and fine-tuned weights) [chitale2023task], allows for selection of higher rank during training without any overhead during inference. Task arithmetic assumes that fine-tuning a pre-trained model pushes the weights away towards a semantically relevant direction in the weight space, thus enabling model compression using simple arithmetic operations with task vectors [chitale2023task]. However, task arithmetic suffers from number of shortcomings including finding optimal scaling coefficients and suboptimal performance in the absence of a memory buffer [zhou2024metagpt].
@@
Typically, when using arbitrary weights, a higher rank usually facilitates learning new tasks but tends to increase forgetting with significant increase in number of parameters, while lower-rank mitigates forgetting with limited adaptation [lu2024adaptive, biderman2024lora]. There often exists a balance between plasticity and stability at a moderately small (but not too small) rank, which may enhance the benefits of CL [lu2024adaptive]. To this end, we introduce PEARL (Parameter-Efficient Adaptive Ranking for lifelong learning with LoRA), a rehearsal-free, parameter-isolation-based PEFT vision CL framework that mitigates catastrophic forgetting through dynamic low-rank adaptation. Specifically, PEARL leverages the information contained in the reference task weights  [^1] We use reference task interchangeably with pre-training when available. Otherwise, it refers to the first task in CL. and dynamically determines the rank of task-specific, layer-specific LoRA components during training based on the current task’s proximity to reference task weights in parameter space. With our dynamic threshold criterion, a close proximity would result in a lower rank allocation and maximum information re-use while a distant proximity results in a larger rank selection, thereby finding an optimal balance between learning and forgetting. To demonstrate the versatility of PEARL, we perform extensive analyses across various datasets, multiple CL scenarios and different architectures, an overview of which can be found in Appendix A.1. Our results demonstrate that PEARL variants consistently outperform the evaluated baselines in all scenarios by a huge margin. Our contributions are as follows:
@@
  • We conduct extensive experiments across 3 different architectures (ResNet, Separable Convolutional Network, and Vision Transformer), 8 distinct settings, 2 CL scenarios, and 2 CL metrics to clearly demonstrate that PEARL significantly outperforms the considered baselines. To further substantiate our claims, we provide 10+ comprehensive analysis in Appendix B.
@@ 2 Related Works
Exemplar and Non-Exemplar Based CL: Experience Rehearsal (ER), which stores and replays samples from previous tasks, has remained a leading approach to addressing the issue of catastrophic forgetting in CL. Several works build upon ER: DER++ [buzzega2020dark] proposed a consistency regularization technique to enhance stability and generalization. Guided by the GWT, TAMiL [bhat2022task] and AGILE [bhat2024mitigating] incorporate information reuse and task-specific, parameter-efficient adapters that re-project the backbone features onto a global workspace. However, these approaches face additional computational and memory overhead due to model surrogates and experience rehearsal. Since storing samples from previous tasks is not always feasible, non-exemplar-based approaches are gaining traction. Non-Exemplar Class-IL (NECIL) is particularly important in scenarios where data confidentiality is crucial due to privacy or security concerns, and where data storage has a limited lifespan. Early contributions to this domain include LwF [li2017learning] and EWC [kirkpatrick2017overcoming], both of which employ regularization techniques to reduce catastrophic forgetting. More recent methods, such as PASS [zhu2021prototype] and IL2A [zhu2021class], have further advanced NECIL by generating prototypes for previous classes without the need to retain the original images. SSRE [zhu2022self] introduced a reparameterization method to balance old and new knowledge, along with additional self-training that utilizes external data as an alternative to exemplars. FeTrIL [petit2023fetril] proposed a framework that combines a fixed feature extractor with a pseudo-feature generator using geometric transformations to achieve a better stability-plasticity balance. FeCAM [goswami2024fecam] addressed the limitations of Euclidean distance when feature distributions are non-stationary and suggested using anisotropic Mahalanobis distance as a substitute for Euclidean distance when employing a frozen feature extractor for prototype classification. Overall, the NECIL benchmark presents a significant challenge as it prohibits the storage of any samples from previous tasks while requiring effective performance in a Class-IL setting. PEARL also aligns with the NECIL benchmark as it operates without any experience rehearsal.
@@
Parameter Efficient Fine-Tuning (PEFT) techniques that have emerged in the space of Large Language Models have been quite pivotal in optimizing model performance with little memory and computational overhead. Seminal works such as Adapter modules [houlsby2019parameter], Prompt Tuning [brown2020language], BitFit [zaken2021bitfit], LoRA [hu2021lora], and DoRA [liu2024dora] have demonstrated the efficacy of selective fine-tuning all the while preserving model generalization and enhancing model adaptability. Specifically, LoRA introduces two compact, trainable low rank matrices into each layer’s weight matrix, allowing model to adapt to new tasks with minimal memory and computational footprint. Within CL, several approaches leverage PEFT techniques to mitigate catastrophic forgetting: Prompt-based methods such as L2P [wang2022learning] and DualPrompt [wang2022dualprompt] introduce task-specific prompts to facilitate learning new tasks while preserving consolidated knowledge. S-Prompt [wang2022s] and CODA-Prompt [smith2023coda] extend these works by employing structural prompts to map discriminative relationships and through Schmidt orthogonalization respectively. In order to maintain an optimal balance between computational complexity and accuracy, adapter-based methods such as SEMA [SEMA] and DIA [DIA] propose self-expanding, dynamic integration of task-specific adapters. Similarly, EASE [zhou2024expandable] creates task-specific sub-spaces by employing lightweight adapter modules for each new task, thereby maintaining resource efficiency. In contrast to adapter-based methods, InfLoRA [liang2024inflora] inserts a small number of LoRA parameters to re-parameterize the pre-trained weights and shows that fine-tuning these inserted parameters is equivalent to fine-tuning the pre-trained weights within a subspace. C-LoRA [DBLP:journals/corr/abs-2304-06027] incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. While these approaches have shown relative success, their implementation is tailored specifically to pre-trained vision transformers and do not consider task dynamics, limiting the potential for information reuse and dynamic resource allocation.
@@
Task Arithmetic [ilharcoediting], centered around task vectors, offers a new paradigm for steering the behavior of pre-trained / reference network towards multiple downstream tasks using task similarity. Inspired by recent works on weight interpolation [frankle2020linear, ilharco2022patching, ainsworth2022git], task vectors are defined to be the element-wise difference between pre-trained / reference network weights before and after fine-tuning on a particular task. A task vector specifies a direction in the weight space of a pre-trained model, such that movement in that direction improves performance on a given task [ilharcoediting]. As these task vectors can be modified and combined together in simple arithmetic operations, new tasks can be learnt via addition while interfering tasks can be forgotten via negation. Inspired by these findings, chitale2023task proposed an approach to continually train transformer-based vision models using low-rank adaptation and task arithmetic. Although task arithmetic is quite lucrative for its simplicity and intuitiveness, it suffers from several shortcomings such as finding optimal scaling coefficients and sub-optimal performance in the absence of memory buffer to finetune the network after weight aggregation [zhou2024metagpt].
@@ 3.2 Information re-use and dynamic resource allocation during CL
As $A$ and $B$ are initialized from SV-Decomposition matrices, at times, the CL model often struggles to find an optimum in the loss landscape resulting in a sub-optimal performance. Therefore, before any further fine-tuning on the current task, we re-initialize the weights of $A$, $B$ along with task-specific normalization (if any) and classification layers. An ablation study on the same can be found in Section 4.
@@ 3.3 Putting it all together
The PEARL framework requires reference task weights to facilitate information re-use across CL tasks. When pre-training is available (Algorithm 1), the pre-trained weights serve as reference task weights. In the absence of pre-training, the first task in CL is considered the reference task (Algorithm 2). For each subsequent CL task, a copy of the reference task weights is fine-tuned for a specified number of epochs on the current task data using cross-entropy loss. We compute the corresponding layer-specific task vector as the element-wise difference between the fine-tuned weights and the reference task weights. Depending on the backbone architecture, the task vectors are computed for convolutional, point-wise convolutional, or linear projections for the key in self-attention, as applicable. The task vector is then decomposed using SV-Decomposition through Equation 2. Since task vectors entail additional information beyond the reference task weights, we model them using dynamic LoRA components. Equations 3 and 4 together define a dynamic rank specific to the current task and layer under consideration. Given the rank $k$, matrices $B$ and $A$ are initialized as described in Equation 5. At times, this initialization may lead to local minima, resulting in suboptimal performance. Therefore, we re-initialize the LoRA matrices, along with any task-specific normalization and classification layers, before further fine-tuning on the current task using cross-entropy loss. This process is repeated for all target layers across the backbone network. After fine-tuning, these task-specific parameters are kept frozen.
@@
Our CL model is trained simultaneously on both Class-IL and Task-IL settings, as their training regimes are identical. During inference in the Task-IL setting, $\theta_{t}=\{f_{\theta^{r}},f_{\theta^{t}},g_{\theta^{t}}\}$ is selected based on the given task-id $t$, and its output is inferred for maximum activation. As task-id is not available in Class-IL inference, each test image passes through every sub-network $\theta_{t}$, and their respective classifier ($g_{\theta^{t}}$) outputs are concatenated and then inferred for maximum activation. For more details on training and inference in Class-IL and Task-IL settings, refer Sections C.2, and C.3.
@@
Table 1: Comparison of Class-IL performance of exemplar-based and NECIL approaches across different datasets and number of tasks. Baselines use randomly initialized ResNet-18 as a backbone. Exemplar-based approaches use a buffer size of 200. NECIL results are taken from [goswami2024resurrecting], while for the rest, we report an average of 3 runs. Best results are marked in bold while second best are underlined.
@@
DER++ [buzzega2020dark] NeurIPS 2020 28.8 47.89 25.42 43.94 17.75 35.89 10.15 27.26
TAMiL [bhat2022task] ICLR 2023 41.43 58.09 32.20 49.87 26.32 43.91 20.46 40.94
AGILE [bhat2024mitigating] CoLLAs 2024 45.73 57.66 33.23 42.88 20.14 43.79 20.19 43.81
LwF [li2017learning] ECCV 2016 45.35 61.94 26.14 46.14 38.81 49.70 27.42 38.77
IL2A [zhu2021class] NeurIPS 2021 - - 31.70 48.40 - - 25.30 42.01
PASS [zhu2021prototype] CVPR 2021 49.75 63.39 37.78 52.18 36.44 48.64 26.58 38.65
SSRE [zhu2022self] CVPR 2022 42.39 56.57 29.44 44.38 30.13 43.20 22.48 34.93
FeTrIL [petit2023fetril] WACV 2023 45.11 60.42 36.69 52.11 29.91 43.99 23.88 36.35
FeCAM [goswami2024fecam] NeurIPS 2024 47.28 61.37 33.82 48.58 25.62 39.85 23.21 35.32
@@
4 Results
Experimental settings: We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL, a complete overview of which can be found in Appendix A.1. To facilitate this evaluation, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE), (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM), (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE), and (iv) PEFT approaches for CL (e.g., L2P, C-LoRA, InfLoRA, SEMA, ADA). We also provide a justification for the choice of these baseline in Appendix A.2. The first three categories utilize randomly initialized ResNet-18 as their backbone, whereas PEFT approaches leverage the ViT-Base-16 model pre-trained on ImageNet-21k. It is important to note that pre-training is often a luxury for smaller, custom models. Consequently, in accordance with the training protocols of the baseline methods, we employ randomly initialized convolutional models, while the ViT model is pre-trained. To demonstrate the versatility of PEARL across different architectures, we explore three distinct backbones: (i) Convolutional networks such as ResNet-18 (R18) and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [li2022blueprint], and (ii) the Vision Transformer (ViT) Vit-Base-16 pre-trained on ImageNet-21k. For each subsequent task, we adapt convolutional layer, point-wise convolutional layer and linear layer projecting key in self-attention using dynamic low-rank adaptation in each of these backbones respectively. More information on backbones and their settings can be found in Appendix C.
Comparison with convolutional CL approaches: Table 1 presents a comparison of PEARL against exemplar-based and NECIL approaches across four different Class-IL settings. We report results for both of our convolutional variants. Since the backbone is randomly initialized without pre-training, PEARL convolutional variants use the first CL task as the reference task and assign LoRA components for subsequent tasks based on their proximity to the first task in the parameter space (Algorithm 2). As shown, PEARL (R18) outperforms all considered baselines without relying on experience rehearsal while exhibiting modest model growth with an increasing number of tasks (see Table 11). Although a larger buffer size might enhance the performance of rehearsal-based approaches, any improvements would likely be outweighed by the increased computational and memory costs [TriRE2023]. Additionally, NECIL approaches do not always re-use information across tasks, which often leads to subpar performance. In contrast, PEARL capitalizes on the information contained in the first task for every subsequent task, achieving an optimal balance between performance and model size. While the absence of a pre-trained model may limit the generalizability of the first task, the framework’s ability to reuse information and dynamically allocate resources allows PEARL to efficiently assimilate information from both the first and current tasks, resulting in superior performance across tasks.

### v4

@@ 1 Introduction
Learning forms the foundation for intelligent systems to adapt to dynamic environments in response to external changes. Unlike conventional deep neural networks (DNNs) that are designed for static data distribution, Continual Learning (CL) is characterized by learning from sequential, dynamic data distribution (Wang et al., 2023b). However, CL has remained a significant challenge for DNNs due to the issue of catastrophic forgetting: A phenomenon in which adaptation to new tasks results in significant deterioration of consolidated knowledge (Masana et al., 2022; De Lange et al., 2021; Ditzler et al., 2015; Robins, 1995). A number of approaches such as rehearsal-based (e.g., Buzzega et al. (2020)), weight-regularization based (e.g., Kirkpatrick et al. (2017)) and parameter-isolation based (e.g., Rusu et al. (2016)) have been proposed in the literature to address the problem of catastrophic forgetting in CL. Although CL is primarily intended to ensure the resource efficiency of model updates (Wang et al., 2023b), the reduction in catastrophic forgetting in these approaches is reliant on increases in computational and memory footprint due to model surrogates and memory buffer. Extrapolating forward this reliance in real world scenarios is both economically and environmentally infeasible.
@@
The increased availability of pre-trained models that demonstrate strong generalization capabilities has led researchers to discover early exploratory methods (Douillard et al., 2022; Jeeveswaran et al., 2023; Gao et al., 2024; Yang et al., 2024), facilitating more efficient integration of new knowledge in CL systems. However, continuous fine-tuning on multiple downstream tasks compromises pre-trained model’s generalization capabilities (Liu et al., 2024a). Recently, Parameter Efficient Fine-Tuning (PEFT) CL techniques (Liang & Li, 2024; Zhou et al., 2024b) are on the rise as they provide lightweight training schedule while avoiding degradation of consolidated knowledge in pre-trained models. However, it is not clear how one can select the size of their rank (in case of LoRA (Hu et al., 2021)) while their performance is very sensitive to rank selection (Valipour et al., 2022). Although choosing a higher rank usually results in optimal performance, it significantly increases the number of learnable parameters in longer task sequences when using large pre-trained models. Task arithmetic (Ilharco et al., 2023), which enables model compression using simple arithmetic operations with task vectors (difference between pre-trained and fine-tuned weights) (Chitale et al., 2023), allows for selection of higher rank during training without any overhead during inference. Task arithmetic assumes that fine-tuning a pre-trained model pushes the weights away towards a semantically relevant direction in the weight space, thus enabling model compression using simple arithmetic operations with task vectors (Chitale et al., 2023). However, task arithmetic suffers from number of shortcomings including finding optimal scaling coefficients and suboptimal performance in the absence of a memory buffer (Zhou et al., 2024c).
@@
Typically, when using arbitrary weights, a higher rank usually facilitates learning new tasks but tends to increase forgetting with significant increase in number of parameters, while lower-rank mitigates forgetting with limited adaptation (Lu et al., 2024; Biderman et al., 2024). There often exists a balance between plasticity and stability at a moderately small (but not too small) rank, which may enhance the benefits of CL (Lu et al., 2024). To this end, we introduce PEARL (Parameter-Efficient Adaptive Ranking for lifelong learning with LoRA), a rehearsal-free, parameter-isolation-based PEFT vision CL framework that mitigates catastrophic forgetting through dynamic low-rank adaptation. Specifically, PEARL leverages the information contained in the reference task weights  [^1] We use reference task interchangeably with pre-training when available. Otherwise, it refers to the first task in CL. and dynamically determines the rank of task-specific, layer-specific LoRA components during training based on the current task’s proximity to reference task weights in parameter space. With our dynamic threshold criterion, a close proximity would result in a lower rank allocation and maximum information re-use while a distant proximity results in a larger rank selection, thereby finding an optimal balance between learning and forgetting. To demonstrate the versatility of PEARL, we perform extensive analyses across various datasets, multiple CL scenarios and different architectures, an overview of which can be found in Appendix B.1. Our results demonstrate that PEARL variants consistently outperform the evaluated baselines in all scenarios by a huge margin. Our contributions are as follows:
@@
  • We conduct extensive experiments across 3 different architectures (ResNet, Separable Convolutional Network, and Vision Transformer), 8 distinct settings, 2 CL scenarios, and 2 CL metrics to clearly demonstrate that PEARL significantly outperforms the considered baselines. To further substantiate our claims, we provide 10+ comprehensive analysis in Appendix C.
@@ 2 Related Works
Exemplar and Non-Exemplar Based CL: Experience Rehearsal (ER), which stores and replays samples from previous tasks, has remained a leading approach to addressing the issue of catastrophic forgetting in CL. Several works build upon ER: DER++ (Buzzega et al., 2020) proposed a consistency regularization technique to enhance stability and generalization. Guided by the GWT, TAMiL (Bhat et al., 2023) and AGILE (Bhat et al., 2024) incorporate information reuse and task-specific, parameter-efficient adapters that re-project the backbone features onto a global workspace. However, these approaches face additional computational and memory overhead due to model surrogates and experience rehearsal. Since storing samples from previous tasks is not always feasible, non-exemplar-based approaches are gaining traction. Non-Exemplar Class-IL (NECIL) is particularly important in scenarios where data confidentiality is crucial due to privacy or security concerns, and where data storage has a limited lifespan. Early contributions to this domain include LwF (Li & Hoiem, 2017) and EWC (Kirkpatrick et al., 2017), both of which employ regularization techniques to reduce catastrophic forgetting. More recent methods, such as PASS (Zhu et al., 2021b) and IL2A (Zhu et al., 2021a), have further advanced NECIL by generating prototypes for previous classes without the need to retain the original images. SSRE (Zhu et al., 2022) introduced a reparameterization method to balance old and new knowledge, along with additional self-training that utilizes external data as an alternative to exemplars. FeTrIL (Petit et al., 2023) proposed a framework that combines a fixed feature extractor with a pseudo-feature generator using geometric transformations to achieve a better stability-plasticity balance. FeCAM (Goswami et al., 2024a) addressed the limitations of Euclidean distance when feature distributions are non-stationary and suggested using anisotropic Mahalanobis distance as a substitute for Euclidean distance when employing a frozen feature extractor for prototype classification. Overall, the NECIL benchmark presents a significant challenge as it prohibits the storage of any samples from previous tasks while requiring effective performance in a Class-IL setting. PEARL also aligns with the NECIL benchmark as it operates without any experience rehearsal.
@@
Parameter Efficient Fine-Tuning (PEFT) techniques that have emerged in the space of Large Language Models have been quite pivotal in optimizing model performance with little memory and computational overhead. Seminal works such as Adapter modules (Houlsby et al., 2019), Prompt Tuning (Brown, 2020), BitFit (Zaken et al., 2021), LoRA (Hu et al., 2021), and DoRA (Liu et al., 2024b) have demonstrated the efficacy of selective fine-tuning all the while preserving model generalization and enhancing model adaptability. Specifically, LoRA introduces two compact, trainable low rank matrices into each layer’s weight matrix, allowing model to adapt to new tasks with minimal memory and computational footprint. Within CL, several approaches leverage PEFT techniques to mitigate catastrophic forgetting: Prompt-based methods such as L2P (Wang et al., 2022e) and DualPrompt (Wang et al., 2022c) introduce task-specific prompts to facilitate learning new tasks while preserving consolidated knowledge. S-Prompt (Wang et al., 2022a) and CODA-Prompt (Smith et al., 2023b) extend these works by employing structural prompts to map discriminative relationships and through Schmidt orthogonalization respectively. In order to maintain an optimal balance between computational complexity and accuracy, adapter-based methods such as SEMA (Wang et al., 2025) and DIA (Li et al., 2025) propose self-expanding, dynamic integration of task-specific adapters. Similarly, EASE (Zhou et al., 2024b) creates task-specific sub-spaces by employing lightweight adapter modules for each new task, thereby maintaining resource efficiency. In contrast to adapter-based methods, InfLoRA (Liang & Li, 2024) inserts a small number of LoRA parameters to re-parameterize the pre-trained weights and shows that fine-tuning these inserted parameters is equivalent to fine-tuning the pre-trained weights within a subspace. C-LoRA (Smith et al., 2023a) incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. While these approaches have shown relative success, their implementation is tailored specifically to pre-trained vision transformers and do not consider task dynamics, limiting the potential for information reuse and dynamic resource allocation.
@@
Task Arithmetic (Ilharco et al., 2023), centered around task vectors, offers a new paradigm for steering the behavior of pre-trained / reference network towards multiple downstream tasks using task similarity. Inspired by recent works on weight interpolation (Frankle et al., 2020; Ilharco et al., 2022; Ainsworth et al., 2022), task vectors are defined to be the element-wise difference between pre-trained / reference network weights before and after fine-tuning on a particular task. A task vector specifies a direction in the weight space of a pre-trained model, such that movement in that direction improves performance on a given task (Ilharco et al., 2023). As these task vectors can be modified and combined together in simple arithmetic operations, new tasks can be learnt via addition while interfering tasks can be forgotten via negation. Inspired by these findings, Chitale et al. (2023) proposed an approach to continually train transformer-based vision models using low-rank adaptation and task arithmetic. Although task arithmetic is quite lucrative for its simplicity and intuitiveness, it suffers from several shortcomings such as finding optimal scaling coefficients and sub-optimal performance in the absence of memory buffer to finetune the network after weight aggregation (Zhou et al., 2024c).
@@ 3.2 Information re-use and dynamic resource allocation during CL
As $A$ and $B$ are initialized from SV-Decomposition matrices, at times, the CL model often struggles to find an optimum in the loss landscape resulting in a sub-optimal performance. Therefore, before any further fine-tuning on the current task, we re-initialize the weights of $A$, $B$ along with task-specific normalization (if any) and classification layers. An ablation study on the same can be found in Section 4.5.
@@ 3.3 Putting it all together
The PEARL framework requires reference task weights to facilitate information re-use across CL tasks. When pre-training is available (Algorithm 1), the pre-trained weights serve as reference task weights. In the absence of pre-training, the first task in CL is considered the reference task (Algorithm 2). For each subsequent CL task, a copy of the reference task weights is fine-tuned for a specified number of epochs ($E_{1}$) on the current task data using cross-entropy loss. We compute the corresponding layer-specific task vector as the element-wise difference between the fine-tuned weights and the reference task weights. Depending on the backbone architecture, the task vectors are computed for convolutional, point-wise convolutional, or linear projections for the key in self-attention, as applicable. The task vector is then decomposed using SV-Decomposition through Equation 2. Since task vectors entail additional information beyond the reference task weights, we model them using dynamic LoRA components. Equations 3 and 4 together define a dynamic rank specific to the current task and layer under consideration. Given the rank $k$, matrices $B$ and $A$ are initialized as described in Equation 5. At times, this initialization may lead to local minima, resulting in suboptimal performance. Therefore, we re-initialize the LoRA matrices, along with any task-specific normalization and classification layers, before further fine-tuning on the current task using cross-entropy loss for another $E_{2}$ epochs. This process is repeated for all target layers across the backbone network. After fine-tuning, these task-specific parameters are kept frozen.
@@
Our CL model is trained simultaneously on both Class-IL and Task-IL settings, as their training regimes are identical. During inference in the Task-IL setting, $\theta_{t}=\{f_{\theta^{r}},f_{\theta^{t}},g_{\theta^{t}}\}$ is selected based on the given task-id $t$, and its output is inferred for maximum activation. As task-id is not available in Class-IL inference, each test image passes through every sub-network $\theta_{t}$, and their respective classifier ($g_{\theta^{t}}$) outputs are concatenated and then inferred for maximum activation. For more details on training and inference in Class-IL and Task-IL settings, refer Sections D.2, and D.3.
4 Results
Experimental settings: We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL, a complete overview of which can be found in Appendix B.1. To facilitate this evaluation, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE), (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM), (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE), and (iv) PEFT approaches for CL (e.g., L2P, C-LoRA, InfLoRA, SEMA, ADA). We also provide a justification for the choice of these baseline in Appendix B.2. The first three categories utilize randomly initialized ResNet-18 as their backbone, whereas PEFT approaches leverage the ViT-Base-16 model pre-trained on ImageNet-21k. It is important to note that pre-training is often a luxury for smaller, custom models. Consequently, in accordance with the training protocols of the baseline methods, we employ randomly initialized convolutional models, while the ViT model is pre-trained. To demonstrate the versatility of PEARL across different architectures, we explore three distinct backbones: (i) Convolutional networks such as ResNet-18 (R18) and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) (Li et al., 2022b), and (ii) the Vision Transformer (ViT) Vit-Base-16 pre-trained on ImageNet-21k. For each subsequent task, we adapt convolutional layer, point-wise convolutional layer and linear layer projecting key in self-attention using dynamic low-rank adaptation in each of these backbones respectively. More information on backbones and their settings can be found in Appendix D.
@@ 4 Results
Table 1: Comparison of Class-IL performance of exemplar-based and NECIL approaches across different datasets and number of tasks. Baselines use randomly initialized ResNet-18 as a backbone. Exemplar-based approaches use a buffer size of 200. NECIL results are taken from Goswami et al. (2024b), while for the rest, we report an average of 3 runs. Best results are marked in bold while second best are underlined.
@@
DER++ (Buzzega et al., 2020) NeurIPS 2020 28.8 47.89 25.42 43.94 17.75 35.89 10.15 27.26
TAMiL (Bhat et al., 2023) ICLR 2023 41.43 58.09 32.20 49.87 26.32 43.91 20.46 40.94
AGILE (Bhat et al., 2024) CoLLAs 2024 45.73 57.66 33.23 42.88 20.14 43.79 20.19 43.81
LwF (Li & Hoiem, 2017) ECCV 2016 45.35 61.94 26.14 46.14 38.81 49.70 27.42 38.77
IL2A (Zhu et al., 2021a) NeurIPS 2021 - - 31.70 48.40 - - 25.30 42.01
PASS (Zhu et al., 2021b) CVPR 2021 49.75 63.39 37.78 52.18 36.44 48.64 26.58 38.65
SSRE (Zhu et al., 2022) CVPR 2022 42.39 56.57 29.44 44.38 30.13 43.20 22.48 34.93
FeTrIL (Petit et al., 2023) WACV 2023 45.11 60.42 36.69 52.11 29.91 43.99 23.88 36.35
FeCAM (Goswami et al., 2024a) NeurIPS 2024 47.28 61.37 33.82 48.58 25.62 39.85 23.21 35.32
@@
4.1 Comparison with convolutional CL approaches
Table 1 presents a comparison of PEARL against exemplar-based and NECIL approaches across four different Class-IL settings. We report results for both of our convolutional variants. Since the backbone is randomly initialized without pre-training, PEARL convolutional variants use the first CL task as the reference task and assign LoRA components for subsequent tasks based on their proximity to the first task in the parameter space (Algorithm 2). As shown, PEARL (R18) outperforms all considered baselines without relying on experience rehearsal while exhibiting modest model growth with an increasing number of tasks (see Table 8). Although a larger buffer size might enhance the performance of rehearsal-based approaches, any improvements would likely be outweighed by the increased computational and memory costs (Vijayan et al., 2023b). Additionally, NECIL approaches do not always re-use information across tasks, which often leads to subpar performance. In contrast, PEARL capitalizes on the information contained in the first task for every subsequent task, achieving an optimal balance between performance and model size. While the absence of a pre-trained model may limit the generalizability of the first task, the framework’s ability to reuse information and dynamically allocate resources allows PEARL to efficiently assimilate information from both the first and current tasks, resulting in superior performance across tasks.